OpenStudy (anonymous):

The volume of a rectangular pyramid is one third the product of the area of its base and its height. Find an expression for the volume of a rectangular pyramid whose base has an area of 3x to the power of 2 + 12x + 9 square feet and whose height is x + 3 feet.

OpenStudy (anonymous):

(1/3)(3x^2 + 12x + 9)(x + 3) = (x^2 + 4x + 3)(x + 3) = (x^3 + 7x^2 + 15x + 9)

OpenStudy (anonymous):

*nod nod* Thank you very much!

OpenStudy (anonymous):

haha your very welcome